### v1

# Lepton seesaw model in a modular $A_{4}$ symmetry

Abstract
We propose a lepton seesaw model to get the same mass-origin of charged-leptons and neutrinos introducing a modular $A_{4}$ symmetry. In this scenario, the mass matrix of charged-leptons is induced via seesaw-like mechanism while the mass matrix of neutrino is realized via an inverse seesaw mechanism with help of vector-like fermions and an isospin singlet and triplet scalar field with nonzero vacuum expectation values. In analyzing the model we concentrate on fixed points of modulus $\tau$ that is favored by a flux compactification of Type IIB string theory. We search for allowed region at nearby these fixed points and find good solutions at nearby $\tau=\omega$ with several good predictions.
I Introduction
Even though charged-leptons and the neutrinos are in the same component in the left-handed $SU(2)$ doublet of the standard model (SM), their masses are often discussed with different manners. This is because the masses of neutrinos are zero in the SM and we need to extend the SM so that we have (tiny but) non-zero neutrino masses. One of the simplest ways is to introduce heavy right-handed neutrinos that is analogy of the other three sectors, and the small neutrino masses can simply realized via seesaw mechanism Yanagida (1979); Minkowski (1977); Mohapatra and Senjanovic (1980); Zee (1980). But in this scenario, there is no relation between the mass matrix of neutrinos and charged-leptons. Thus it is interesting if we find a unified description of generating masses of them.
In this paper, we study a possibility of the same origin of the charged-leptons and the neutrino mass matrices Lee et al. (2021), which is called ”lepton seesaw”, introducing several exotic fermions (vector-like fermions) and scalar fields (an isospin triplet scalar of hyper-charge 1 and singlet scalar of hyper-charge zero with nonzero vacuum expectations). In order to realize our model, we impose a modular flavor $A_{4}$ group that has a strong symmetry to get our desired terms as well as restricting some forms of mass matrices. Since 2017 Feruglio (2019), this symmetry, especially modular $A_{4}$, has widely been applied to various phenomenologies in flavor physics Criado and Feruglio (2018); Kobayashi et al. (2018); Okada and Tanimoto (2019); Nomura and Okada (2019); Okada and Tanimoto (2021a); de Anda et al. (2020); Novichkov et al. (2019); Nomura and Okada (2021); Okada and Orikasa (2019); Ding et al. (2019a); Nomura et al. (2020); Kobayashi et al. (2019); Asaka et al. (2020a); Zhang (2020); Ding et al. (2019b); Kobayashi et al. (2020); Nomura et al. (2021); Wang (2020); Okada and Shoji (2020); Okada and Tanimoto (2023); Behera et al. (2022a, b); Nomura and Okada (2022, 2020); Asaka et al. (2020b); Okada and Tanimoto (2021b); Nagao and Okada (2022); Okada and Tanimoto (2021c); Kang et al. (2022); Ding et al. (2024); Ding and King (2023); Nomura et al. (2024); Kobayashi et al. (2024); Petcov and Tanimoto (2024); Kobayashi and Tanimoto (2023); Nomura and Okada (2024); Qu and Ding (2024). In the model of this work, modular $A_{4}$ symmetry plays two important roles; (1) Forbidding unnecessary terms to realize lepton seesaw, (2) Providing predictions constraining structure of Yukawa couplings.
After constructing the model we perform numerical analysis to find some predictions in neutrino sector imposing experimental constraints. In our chi square numerical analysis, we concentrate on some specific region at nearby fixed points of modulus $\tau$ that is a new degree of freedom via modular symmetry. This fixed points are favored by a flux compactification of Type IIB string theory Ishiguro et al. (2021) and we find allowed region at nearby $\tau=\omega$ that reserves a remnant $Z_{3}$ symmetry, and obtain several predictions such as phases and masses on neutrino sector.
This paper is organized as follows. In Sec. II, we review our setup of the modular $A_{4}$ assignments introducing new particles and explain how to realize our scenario by showing the charged-fermion sector and the neutral fermion sector. In Sec. III, we demonstrate chi-square numerical analyses concentrating on fixed points and show what kinds of predictions we obtain. Finally, we summarize and conclude in Sec. IV.
II Model setup
 $L_{L}$  $\overline{\ell_{R}}$  $\overline{L^{\prime}_{R}}$  ${L^{\prime}_{L}}$  $H$   $\Delta$   $\varphi$
$SU(2)_{L}$ $\bm{2}$ $\bm{1}$ $\bm{2}$ $\bm{2}$ $\bm{2}$ $\bm{3}$ $\bm{1}$
$U(1)_{Y}$ $-\frac{1}{2}$ $1$ $\frac{1}{2}$ $-\frac{1}{2}$ $\frac{1}{2}$ $1$ 00
$A_{4}$ $\bm{3}$$\{\bar{\bm{1}}\}$$\{\bar{\bm{1}}\}$$\{\bm{1}\}$ $\bm{1}$ $\bm{1}$ $\bm{1}$
$-k_{I}$ $+3$ $-2$ $-2$ $+2$ 00 $-4$ $-3$
Table 1: Charge assignments of the fermions and bosons under $SU(2)_{L}\otimes U(1)_{Y}\otimes A_{4}$ where $-k_{I}$ is the number of modular weight. Here, $\{\bm{1}\}=\{1,1^{\prime},1^{\prime\prime}\}$ indicates assignment of $A_{4}$ singlets.
Here, we review our model. At first, we introduce three isospin doublet vector-like leptons $L^{\prime}(\equiv[N,E]^{T}$) with $U(1)_{Y}=-1/2$, an isospin triplet scalar field $\Delta$ with $U(1)_{Y}=1$, and an isospin singlet scalar field $\varphi(\equiv(v^{\prime}+r)/\sqrt{2})$ with $U(1)_{Y}=0$, in addition to the SM fields. The SM Higgs is denoted by $H$ and its vacuum expectation value (VEV) is defined by $\langle H\rangle=[0,v/\sqrt{2}]^{T}$. We also require $\varphi$ and $\Delta$ to develop VEVs denoted by $v^{\prime}$ and $v_{\Delta}$ respectively. The SM right-handed leptons $\ell_{R}$, and $L^{\prime}$ are assigned to be $\{\bm{1}\}=\{1,1^{\prime},1^{\prime\prime}\}$ under the $A_{4}$ symmetry, while the left-handed SM leptons $L_{L}\equiv[\nu_{L},\ell_{L}]^{T}$ are assigned to be triplet under the $A_{4}$ symmetry. All the scalar fields are trivial singlet under the $A_{4}$ symmetry. In addition, we impose the modular weight for each field; $+3$ for $L_{L}$, $-2$ for $\overline{\ell_{R}}$, $-2$ for $\overline{L^{\prime}_{R}}$, $+2$ for ${L^{\prime}_{L}}$, 00 for $H$, $-4$ for $\Delta$, and $-3$ for $\varphi$. All fields and their assignments are summarized in TABLE 1. These assignments forbids undesired mass terms at tree-level such as $\overline{\ell_{R}}{\tilde{H}}L_{L}$, $\overline{L^{C}_{L}}(i\tau_{2})\Delta L_{L}$ and $\overline{L^{C}_{L}}(i\tau_{2})\Delta L^{\prime}_{L}$ that breaks our lepton seesaw scenario. These terms are forbidden as the following manner; $\overline{\ell_{R}}{\tilde{H}}L_{L}$ and $\overline{L^{C}_{L}}(i\tau_{2})\Delta L^{\prime}_{L}$ are not allowed since there is no weight $-1$ modular form; $\overline{L^{C}_{L}}(i\tau_{2})\Delta L_{L}$ is forbidden since there is no weight $-2$ modular form.
Under these symmetries, the renormalizable Lagrangian is given by
$-{\cal L}_{\ell}=[\overline{\ell_{R}}{\tilde{H}}L^{\prime}_{L}]+[Y^{(2)}_{3}\overline{L^{\prime}_{R}}L_{L}\varphi]+[\overline{L^{\prime C}_{L}}(i\tau_{2})\Delta L^{\prime}_{L}]+[\overline{L^{\prime}_{R}}(i\tau_{2})\Delta^{\dagger}L^{\prime C}_{R}]+[M_{L^{\prime}}\overline{L^{\prime}_{R}}L^{\prime}_{L}]+{\rm h.c.},$ (1)
where $\tau_{2}$ is the second Pauli matrix, and []delimited-[][$\cdots$][ ⋯ ] represents a trivial $A_{4}$ singlet ${\bm{1}}$ abbreviating the free parameters that are explicitly shown later. $Y^{(2)}_{3}\equiv[y_{1},y_{2},y_{3}]^{T}$ is modular coupling that is uniquely fixed when modulus $\tau$ is determined Feruglio (2019). The Higgs potential is given by
${\cal V}$ $=-\mu_{H}^{2}|H|^{2}-\mu_{\Delta}^{2}{\rm Tr}[\Delta^{\dagger}\Delta]-\mu_{\varphi}|\varphi|^{2}+\mu(H^{T}(i\tau_{2})\Delta^{\dagger}H+{\rm h.c.})$
$+\lambda_{H}|H|^{4}+\lambda_{\Delta}{\rm Tr}[|\Delta^{\dagger}\Delta|^{2}]+\lambda^{\prime}_{\Delta}{\rm det}[\Delta^{\dagger}\Delta]+\lambda_{\varphi}|\varphi|^{4}+\lambda_{H\Delta}|H|^{2}{\rm Tr}[\Delta^{\dagger}\Delta]$
$+\lambda^{\prime}_{H\Delta}\sum_{i}^{3}(H^{\dagger}\tau_{i}H){\rm Tr}[\Delta^{\dagger}\tau^{i}\Delta]{+\lambda_{H\varphi}|H|^{2}|\varphi|^{2}}+\lambda_{\Delta\varphi}{\rm Tr}[\Delta^{\dagger}\Delta]|\varphi|^{2},$ (2)
where $\tau_{i}\ (i=1,2,3)$ are Pauli matrices and some of free parameters includes modular forms; e.g., $\mu_{\Delta}\equiv\mu^{\prime}_{\Delta}/(i\tau-i\tau^{*})^{4}$. The scalar potential is similar to that of Higgs triplet model with a little modification due to one scalar singlet extension. In this work we assume the parameters in the potential satisfy phenomenological constraints and realize VEVs of $H$, $\Delta$ and $\varphi$.
II.1 Charged-lepton mass matrix
The mass matrix of charged-lepton originates from the following terms after spontaneous electroweak symmetry breaking;
$[\overline{\ell_{R}}{\tilde{H}}L^{\prime}_{L}]+[Y^{(2)}_{3}\overline{L^{\prime}_{R}}L_{L}\varphi]+[M_{L^{\prime}}\overline{L^{\prime}_{R}}L^{\prime}_{L}]\supset m\overline{\ell_{R}}E_{L}+m^{\prime}\overline{E_{R}}\ell_{L}+M_{L^{\prime}}\overline{E^{\prime}_{R}}E_{L},$ (3)
where each form of $m$, $m^{\prime}$, and $M_{L^{\prime}}$ is given by
$m=\frac{v}{\sqrt{2}}\left(\begin{array}[]{ccc}a_{e}&0&0\\ 0&b_{e}&0\\ 0&0&c_{e}\end{array}\right),\ m^{\prime}=\frac{v^{\prime}}{\sqrt{2}}\left(\begin{array}[]{ccc}f_{1}&0&0\\ 0&f_{2}&0\\ 0&0&f_{3}\end{array}\right)\left(\begin{array}[]{ccc}y_{1}&y_{3}&y_{2}\\ y_{2}&y_{1}&y_{3}\\ y_{3}&y_{2}&y_{1}\end{array}\right),\ M_{L^{\prime}}=\left(\begin{array}[]{ccc}M_{1}&0&0\\ 0&M_{2}&0\\ 0&0&M_{3}\end{array}\right),$ (16)
where $\{a_{e},b_{e},c_{e}\}$, $\{f_{1},f_{2},f_{3}\}$, and $\{M_{1},M_{2},M_{3}\}$ can be real parameters without loss of generality. Then, in basis of $[\ell_{L},E_{L}]^{T}$, the mass matrix of charged-lepton sector is given by
${\cal M}_{e}=\left(\begin{array}[]{cc}0&m\\ m^{\prime}&M_{L^{\prime}}\end{array}\right)$ (19)
When we assume to be $m,m^{\prime}\ll M_{L^{\prime}}$ among mass matrices, we can treat the charged-lepton mass matrix as a seesaw-like as follows. First of all, ${\cal M}_{e}$ is diagonalized by
$\left(\begin{array}[]{cc}m_{e}&0\\ 0&M_{e}\end{array}\right)=O^{\dagger}_{R}{\cal M}_{e}O_{L}.$ (22)
Due to the mass hierarchies among $m,m^{\prime},M_{L^{\prime}}$, we write the mixing matrix and mass matrices as follows:
$O_{L}\simeq\left(\begin{array}[]{cc}1&-\theta_{L}\\ \theta_{L}&1\end{array}\right),\quad\theta_{L}\approx-M^{-1}m^{\prime},$ (25)
$m_{e}\approx mM_{L^{\prime}}^{-1}m^{\prime},\quad M_{e}\approx M_{L^{\prime}}.$ (26)
Since $M_{L^{\prime}}$ is diagonal, we do not need to diagonalize it more. $m_{e}$ is diagonalized by $D_{e}=V_{e_{R}}m_{e}V_{e_{L}}$. Therefore, $|D_{e}|^{2}=V^{\dagger}_{e_{L}}m^{\dagger}_{e}m_{e}V_{e_{L}}$ where $D_{e}=$ diag.$(m_{e},m_{\mu},m_{\tau})$. Since $m_{e}$ is given by
$\frac{vv^{\prime}}{2}\left(\begin{array}[]{ccc}\frac{a_{e}f_{1}}{M_{1}}&0&0\\ 0&\frac{b_{e}f_{2}}{M_{2}}&0\\ 0&0&\frac{c_{e}f_{3}}{M_{3}}\end{array}\right)\left(\begin{array}[]{ccc}y_{1}&y_{3}&y_{2}\\ y_{2}&y_{1}&y_{3}\\ y_{3}&y_{2}&y_{1}\end{array}\right),$ (33)
$\{\frac{a_{e}f_{1}}{M_{1}},\frac{b_{e}f_{2}}{M_{2}},\frac{c_{e}f_{3}}{M_{3}}\}$ are fixed to fit the three observed charged-lepton masses by the following relations:
${\rm Tr}[m^{\dagger}_{e}m_{e}]=|m_{e}|^{2}+|m_{\mu}|^{2}+|m_{\tau}|^{2},\quad{\rm Det}[m^{\dagger}_{e}m_{e}]=|m_{e}|^{2}|m_{\mu}|^{2}|m_{\tau}|^{2},$
$({\rm Tr}[m^{\dagger}_{e}m_{e})^{2}-{\rm Tr}[(m^{\dagger}_{e}m_{e})^{2}]=2(|m_{e}|^{2}|m_{\mu}|^{2}+|m_{\mu}|^{2}|m_{\tau}|^{2}+|m_{e}|^{2}|m_{\tau}|^{2}).$ (34)
II.2 Neutral fermion mass matrix
The mass matrix of neutral fermions comes from the following terms after spontaneous electroweak symmetry breaking;
$[\overline{L^{\prime C}_{L}}(i\tau_{2})\Delta L^{\prime}_{L}]+[\overline{L^{\prime}_{R}}(i\tau_{2})\Delta^{\dagger}L^{\prime C}_{R}]+[Y^{(2)}_{3}\overline{L^{\prime}_{R}}L_{L}\varphi]+[M_{L^{\prime}}\overline{L^{\prime}_{R}}L^{\prime}_{L}]$
$\supset\delta_{N}\overline{N_{L}^{C}}N_{L}+\delta_{\bar{N}}\overline{N_{R}}N_{R}^{C}+m^{\prime}\overline{E_{R}}\ell_{L}+M_{L^{\prime}}\overline{E^{\prime}_{R}}E_{L},$ (35)
where $\delta_{N}$ is given by
$\delta_{N}=\frac{v_{\Delta}}{\sqrt{2}}\left(\begin{array}[]{ccc}y_{\Delta}&0&0\\ 0&0&y^{\prime}_{\Delta}\\ 0&y^{\prime}_{\Delta}&0\end{array}\right)=\frac{v_{\Delta}y_{\Delta}}{\sqrt{2}}\left(\begin{array}[]{ccc}1&0&0\\ 0&0&\tilde{y}^{\prime}_{\Delta}\\ 0&\tilde{y}^{\prime}_{\Delta}&0\end{array}\right)\equiv\frac{v_{\Delta}y_{\Delta}}{\sqrt{2}}\tilde{\delta}_{N},$ (42)
and $\delta_{\bar{N}}$ is three by three symmetric matrix with complex values. Note here that $\delta_{\bar{N}}$ does not appear in the neutrino mass matrix as can be seen later. Thus, we will not explicitly write it down.
When we impose the following mass hierarchies $\delta_{N}\ll m^{\prime}\ll M_{L^{\prime}}$, one finds the inverse neutrino mass matrix as follows:
$m_{\nu}=(M^{-1}_{L^{\prime}}m^{\prime})^{T}\delta_{N}(M^{-1}_{L^{\prime}}m^{\prime}).$ (43)
As we can see, $M_{L^{\prime}},\ m^{\prime}$ also contributes to the charged-lepton mass matrix. Furthermore, the mass hierarchies $\delta_{N}\ll m^{\prime},M_{L^{\prime}}$ would be appropriate assumption since $\delta_{N}$ is generated via $v_{\Delta}$ that is restricted by a few GeV to satisfy the constraint on the $\rho$ parameter Zyla et al. (2020). The mass eigenvalues for the active neutrinos $D_{\nu}=\{D_{\nu_{1}},D_{\nu_{2}},D_{\nu_{3}}\}$ are obtained by diagonalizing the mass matrix as $D_{\nu}\equiv V_{\nu}^{T}m_{\nu}V_{\nu}$. Therefore, we obtain $V_{\nu}^{\dagger}m_{\nu}^{\dagger}m_{\nu}V_{\nu}={\rm diag.}(|D_{\nu_{1}}|^{2},|D_{\nu_{2}}|^{2},|D_{\nu_{3}}|^{2})$ where $D_{\nu_{1,2,3}}$ are the three neutrino mass eigenvalues. The Pontecorvo-Maki-Nakagawa-Sakata (PMNS) mixing matrix $U(\equiv U_{PMNS})$ is thus defined by $U=V_{L}^{\dagger}V_{\nu}$ where $V_{L}$ comes from diagonalization of charged-lepton mass matrix. For convenience, we here define dimensionless neutrino mass matrix; $m_{\nu}\equiv\kappa\tilde{m}_{\nu}$ and $\kappa\equiv\frac{v_{\Delta}y_{\Delta}}{\sqrt{2}}$ being a flavor independent mass dimensional parameter. Therefore, $\kappa$ can be rewritten in terms of rescaled neutrino mass eigenvalues $\tilde{D}_{\nu}(\equiv D_{\nu}/\kappa)$ and atmospheric neutrino mass-squared difference $\Delta m_{\rm atm}^{2}$ as follows:
$({\rm NH}):\ \kappa^{2}=\frac{|\Delta m_{\rm atm}^{2}|}{\tilde{D}_{\nu_{3}}^{2}-\tilde{D}_{\nu_{1}}^{2}},\quad({\rm IH}):\ \kappa^{2}=\frac{|\Delta m_{\rm atm}^{2}|}{\tilde{D}_{\nu_{2}}^{2}-\tilde{D}_{\nu_{3}}^{2}},$ (44)
where NH and IH respectively stand for the normal and inverted hierarchies. Subsequently, the solar neutrino mass-squared difference is determined by the relation
$\Delta m_{\rm sol}^{2}={\kappa^{2}}({\tilde{D}_{\nu_{2}}^{2}-\tilde{D}_{\nu_{1}}^{2}}).$ (45)
This value should be within the allowed range of the experimental value. Later in numerical analysis, we will adopt NuFit 5.2 Esteban et al. (2020) to experimental ranges of neutrino observables. The effective mass for neutrinoless double beta decay can be written by
$\langle m_{ee}\rangle=\kappa|\tilde{D}_{\nu_{1}}\cos^{2}\theta_{12}\cos^{2}\theta_{13}+\tilde{D}_{\nu_{2}}\sin^{2}\theta_{12}\cos^{2}\theta_{13}e^{i\alpha_{2}}+\tilde{D}_{\nu_{3}}\sin^{2}\theta_{13}e^{i(\alpha_{3}-2\delta_{CP})}|.$ (46)
A predicted value of $\langle m_{ee}\rangle$ is constrained by the current KamLAND-Zen data and could be measured in future Abe et al. (2024). Currently the upper bound is found as $\langle m_{ee}\rangle<(36-156)$ meV at 90 % confidence level where the range of the bound is due to the use of different method in estimating nuclear matrix elements. Furthermore the sum of neutrino masses is constrained by the minimal standard cosmological model $\Lambda$CDM $+\sum D_{\nu_{i}}$ that provides the upper bound $\sum D_{\nu}\leq$ 120 meV Vagnozzi et al. (2017); Aghanim et al. (2020), although it becomes weaker if the data are analyzed in the context of extended cosmological models Olive et al. (2014). Recently, DESI and CMB data combination provides more stringent upper bound on the sum as $\sum D_{\nu}\leq$ 72 meV Adame et al. (2024). These two observable $\langle m_{ee}\rangle$ and $\sum D_{\nu}$ are also discussed as our prediction in the numerical analysis. Direct search for neutrino mass is done by the Karlsruhe Tritium Neutrino (KATRIN) experiment Aker et al. (2022), which is the first sub-eV sensitivity on $m_{\nu_{e}}^{2}=(0.26\pm 0.34)$ eV2 at 90 % CL. Here, $m_{\nu_{e}}^{2}\equiv\sum_{i}|(U_{\rm PMNS})_{ei}|^{2}D^{2}_{\nu_{i}}$. We also take this experiment in account in our numerical analysis below.
III Numerical analysis and phenomenology
In this section, we perform chi square numerical analysis to fit the neutrino data and discuss several features of our model. Furthermore, we concentrate on the region at nearby two representative fixed points $\tau=\{i,\omega\}$ and search for the allowed region within the range of $5\sigma$ with five degrees of freedom. However, we would not find any solutions to satisfy the neutrino oscillation data in case of $\tau=i$. Thus, we focus on discussing features in case of $\tau=\omega$ below.
We randomly select whole our rescaled input parameters within the following ranges $[10^{-3}-10^{3}]$ except for $|\tilde{y}^{\prime}_{\Delta}|$ that is $[10^{-5}-1]$ through our numerical analysis.
Figure 1: Numerical $\Delta\chi^{2}$ analyses in case of NH at nearby $\tau=\omega$, where the blue color represents the range of $0-1$, the green $1-2$, the yellow $2-3$, and the red $3-5$ $\sigma$ standard deviation estimated by $\sqrt{\Delta\chi^{2}}$. The left figure shows the allowed range of $\tau$ where the black solid line is the boundary of the fundamental domain at $|\tau|=1$. The central figure represents the allowed range of Majorana phases. The right figure demonstrates the allowed Dirac CP phase $\delta_{\rm CP}$ in terms of sum of neutrino masses $\sum D_{\nu}$, where the red dotted line is the cosmological upper bound while the pink dotted line is upper bound on the DESI and CMB data.
III.1 NH
In this subsection, we discuss the allowed regions in case of NH. In Figs. 1, the blue color represents the range of $0-1$, the green $1-2$, the yellow $2-3$, and the red $3-5$ $\sigma$ standard deviation estimated by $\sqrt{\Delta\chi^{2}}$. The left figure shows the allowed range of $\tau$ where the black solid line is the boundary of the fundamental domain at $|\tau|=1$. It suggests that we have enough solutions at nearby $\tau=\omega$. The central figure represents the allowed range of Majorana phases $\alpha_{21}$ and $\alpha_{31}$. We find $\alpha_{21}$ is almost zero degree but $\alpha_{31}$ is allowed by the ranges [$177.5^{\circ}-178.5^{\circ}$] and [$181^{\circ}-182.5^{\circ}$].
The right figure of Fig. 1 demonstrates the allowed region of Dirac CP phase $\delta_{\rm CP}$ in terms of sum of neutrino masses $\sum D_{\nu}$, where the red dotted line is the cosmological upper bound while the pink dotted line is upper bound from the combination of the DESI and CMB data. $\delta_{\rm CP}$ is allowed by the ranges of [$80^{\circ}-100^{\circ}$] and [$240^{\circ}-300^{\circ}$]. $\sum D_{\nu}$ is allowed by the range of [$0.22\ {\rm eV}-0.35\ {\rm eV}$].
Figure 2: Numerical $\Delta\chi^{2}$ analyses in case of NH at nearby $\tau=\omega$, where color legends are the same as the ones of Fig. 1. These figures show the allowed range of the neutrinoless double beta decay $\langle m_{ee}\rangle$ in terms of the lightest neutrino mass $D_{\nu_{1}}$ (left), $\sum D_{\nu}$, and $m_{\nu_{e}}^{2}$ (right) in unit of eV.
In Fig. 2, we show some predictions in the case of NH at nearby $\tau=\omega$, where color legends are the same as the ones of Fig. 1. The red and gray dotted lines respectively correspond to upper and lower bounds of $\langle m_{ee}\rangle$ experiment at 90 % confidence level. These figures show the allowed range of the neutrinoless double beta decay $\langle m_{ee}\rangle$ in terms of the lightest neutrino mass $D_{\nu_{1}}$ (left), $\sum D_{\nu}$, and $m_{\nu_{e}}^{2}$ (right) in unit of eV. These suggest each of allowed region as follows: $\langle m_{ee}\rangle=[0.07-0.12]$ [eV] that is completely within the range of the experiment. The lightest neutrino mass goes over the range of $D_{\nu_{1}}=[0.07-0.11]$ [eV] whose range is similar to the one of neutrinoless double beta decay. Even though our $\sum D_{\nu}$ does not satisfy the cosmological bound $0.12$ [eV], $m_{\nu_{e}}^{2}$ is totally safe to the experimental upper bound.
III.2 IH
Figure 3: Numerical $\Delta\chi^{2}$ analyses in case of IH at nearby $\tau=\omega$, where color legends and the captions are the same as the ones of Fig. 1.
In this subsection, we discuss the allowed regions in case of IH. In Figs. 3, we show the same kinds of figures as in Figs. 1, and color legends and the captions are the same as the ones of Fig. 1. The left figure suggests that we have enough solutions at nearby $\tau=\omega$, but points are more localized than the case of NH. The central figure implies that $\alpha_{21}$ is almost zero degree that is the same as the case of NH. We find $\alpha_{31}$ is localized at nearby 180 that is also similar tendency as the case of NH.
The right figure of Fig. 3 demonstrates that $\delta_{\rm CP}$ is allowed by the ranges of [$260^{\circ}-280^{\circ}$] while $\sum D_{\nu}$ is allowed by the range of [$0.44\ {\rm eV}-0.50\ {\rm eV}$] that is a typical tendency.
Figure 4: Numerical $\Delta\chi^{2}$ analyses in case of IH at nearby $\tau=\omega$, where color legends are the same as the ones of Fig. 1. These figures show the allowed range of the neutrinoless double beta decay $\langle m_{ee}\rangle$ in terms of the lightest neutrino mass $D_{\nu_{1}}$ (left), $\sum D_{\nu}$, and $m_{\nu_{e}}^{2}$ (right) in unit of eV.
In Fig. 4, we show the same kinds of figures as in Figs. 2, and color legends and the captions are the same as the ones of Fig. 2. These suggest each of allowed region as follows: $\langle m_{ee}\rangle=[0.14-0.16]$ [eV] that is nearby the range of upper bound of this experiment. Therefore, this case has high testability that could be confirmed/excluded in near future. The lightest neutrino mass goes over the range of $D_{\nu_{3}}=[0.14-0.16]$ [eV] whose range is similar to the one of neutrinoless double beta decay. The tendency is likely to the case of NH. $m_{\nu_{e}}^{2}$ has the same tendency as the case of NH. Therefore, it is totally safe to the experimental upper bound.
IV Summary and discussion
We have proposed a lepton seesaw model in a modular $A_{4}$ symmetry. The charged-lepton mass matrix and the neutrino mass matrix have the same origin associated with vector-like fermions. As a result, the charged-lepton mass matrix is induced via seesaw-like mechanism, while the neutrino mass matrix is generated through the inverse seesaw mechanism where the tiny Majorana mass matrix is supposed to be obtained via the VEV of $\Delta$.
In addition we have searched for allowed regions at nearby fixed points which are favored by a flux compactification of Type IIB string theory Ishiguro et al. (2021), and we have found the solutions at nearby $\tau=\omega$. In our chi square numerical analysis, we have found large value of sum of the neutrino masses which exceed the cosmological upper bound $0.12$ eV for both the cases of NH and IH. However, our solutions are totally safe for the direct search for neutrino mass which is done by the Karlsruhe Tritium Neutrino experiment for both the cases. In addition, our region of neutrinoless double beta decay is localized at nearby the experimental bound, and therefore it would be tested near future. We have obtained narrow ranges of Majorana and CP phases which would also be verifiable via some experiments.
Finally, we would like to mention our next task. Even though we have successfully built our lepton seesaw model which has the same mass origin from vector-like fermions, our model has an assumption among mass hierarchies such as $m^{\prime}\ll M_{L^{\prime}}$. This assumption is in a sense very important to construct the seesaw model. Thus, we will move to a model that would have appropriate and natural mass hierarchies among these mass matrices.
Acknowledgments
The work was supported by the Fundamental Research Funds for the Central Universities (T. N.).